40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 12 00:59:45 2011, ryan, rana, Update, PSL, Return of the PSL temperature box 
    Reply  Sat Jul 30 19:06:25 2011, ryan, rana, Update, PSL, Return of the PSL temperature box rctempbox.png
       Reply  Sat Jul 30 20:41:50 2011, rana, Update, PSL, RefCav Stabilization back on Untitled.png
Message ID: 5071     Entry time: Sat Jul 30 19:06:25 2011     In reply to: 4697     Reply to this: 5072
Author: ryan, rana 
Type: Update 
Category: PSL 
Subject: Return of the PSL temperature box 

Quote:

The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.

 Forgot to do this in May. Have just changed the values in the psl.db file now as well as updating them live via Probe.

To make the appropriate change, I took the measured offset (5.31 deg) and added 2x this to the EGUF and EGUL field for the MINCO_MEAS channel. (see instructions here)

Committed the .db file to the SVN.

attached plot shows 8 days of trend with 5.31 degC added to the black trace using the XMGRACE Data Set Transformations

Attachment 1: rctempbox.png  39 kB  | Hide | Hide all
rctempbox.png
ELOG V3.1.3-